OpenStudy (swag):

What is the slope of the line 5x + 2y = 10?

terenzreignz (terenzreignz):

Any linear equation in two variables can be written in the form y = mx + b This is called the slope-intercept form, and is possibly the commonest way of writing linear equations. Could you try manipulating your equation into this form?

Parth (parthkohli):

Hey @SWAG, do you know how to get this equation in the form \(y = \cdots\)?

Parth (parthkohli):

I know that you have been taught something like "solving for \(y\)".

OpenStudy (swag):

Yes

terenzreignz (terenzreignz):

@ParthKohli it kind of sounds like you've been stalking--- "I know that you..." @SWAG So could you solve for y?

Parth (parthkohli):

So can you solve for y?

OpenStudy (swag):

y = 5-(5x/2)

OpenStudy (anonymous):

@SWAG http://www.algebra.com/algebra/homework/Linear-equations/Linear-equations.faq.question.216340.html that explains your question in depth

Parth (parthkohli):

Yeah, so now look out for the coefficient of \(x\): that's the slope

terenzreignz (terenzreignz):

@SWAG Might be clearer like this: \[\huge y = -\frac{5}{2}x+5\]

OpenStudy (swag):

Thanks Guys

terenzreignz (terenzreignz):

Okay, shortcut time :D Given your equation ax + by = c The slope of the line is ALWAYS \[\huge m=-\frac{a}{b}\]

OpenStudy (swag):

Thanks man.

terenzreignz (terenzreignz):

No problem :)